Lots of activity at the high bay today. Kyle and I continued our alignment of the "instrument" optics and were finished by 2PM and have started finishing up the rest of the instrument (putting on the shielding, cryogenics, and wiring). Jeff removed one of the baffles from inside the cryostat and removed our neutral density filter, used in Palestine to attenuate the signal so our detectors wouldn't be saturated and useless there, in preparation for flying at balloon altitudes.
On the ACS/gondola side, Britt continued rewiring the solar power crates and Chappy continued work on one of the star cameras, preparing to purge the pressure vessel with nitrogen gas. We also brought our rotator in and Michele set about cleaning and re-lubricating the bearings. Jeff had a little bit of time free so he helped Britt by drilling some holes on the gondola for the charge controller covers.
We also had a subset of the Aussie crew stop by to check out what we were doing. And in the evening a few of us went to a screening of one episode of the "Frozen Earth" documentary series. Even though it is Halloween, McMurdo's big Halloween party was last weekend, and I missed it. In the spirit of the day and the location, though, Chappy dressed as a penguin.